OpenStudy (anonymous):

A sports team is building a new stadium on a rectangular lot of land. The lot measures 8x by 12x and the field will be 3x by 6x. How much land will be left for the bleachers, restrooms, and other parts of the stadium? (1 point) 78x2 114x2 96x2 30x2 How do you get the answer?

OpenStudy (johnweldon1993):

So the land is 

OpenStudy (johnweldon1993):

The field will be 

OpenStudy (johnweldon1993):

So we take the area of the entire land..and subtract from that the area of the stadium we are building \[\large \text{Area of Land} = \text{Length} \times \text{Width}\] \[\large \text{Area of Land} = 12x \times 8x\] \[\large \text{Area of Land} = 96x^2\] And \[\large \text{Area of Stadium} = \text{Length} \times \text{Width}\] \[\large \text{Area of Stadium} = 6x \times 3x\] \[\large \text{Area of Stadium} = 18x^2\] So Area left would be \[\large \text{Area Remaining} = 96x^2 - 18x^2 = \space ?\]

OpenStudy (anonymous):

78x^2! Thanks soo much!

OpenStudy (johnweldon1993):

Anytime!